### v1

# Memories of Ian G. Macdonald

Abstract
This is a slightly edited translation of a paper in Dutch which appeared in Nieuw Archief voor Wiskunde (5) 25 (2024), No.2, 87–90 on the occasion of I. G. Macdonald’s death in 2023, and aimed at a very broad mathematical audience. First we review some of Macdonald’s most important older results. Then we focus on the period 1985–1995 when Macdonald often visited the Netherlands and there was much interaction between his work, notably the Macdonald polynomials, and the work by the authors. We end with some glimpses about Macdonald as a person.
Macdonald giving a lecture in Amsterdam on January 11, 2002.
Ian Grant Macdonald was an English mathematician, known for his contributions in algebra, geometry, Lie theory, combinatorics and special functions. He passed away on August 8, 2023 at the age of 94. See obituaries by Peter Cameron [2] and Arun Ram [35]. Ram also wrote a longer memorial paper [36]. Much earlier A. O. Morris wrote a nice biographical paper [29] about him. In the present memorial article we will in particular focus on our contacts with Macdonald starting in the eighties of the previous century.
Macdonald studied at Trinity College in Cambridge, where he graduated in 1952. Instead of subsequently preparing a PhD thesis, he worked during five years as a civil servant. But then he could no longer resist to return to university, where he had consecutive appointments in Manchester, Exeter, Oxford and London. He was elected Fellow of the Royal Society in 1979. In 2002 he received an honorary doctorate of the University of Amsterdam. In his word of thanks he expressed his great satisfaction to finally having received the doctor’s degree.
Below we will first discuss some highlights of Macdonald’s work in the seventies of the previous century. Next we turn to his work starting in the middle of the eighties on orthogonal polynomials associated with root systems. It was in particular this topic where Macdonald’s work had a close relationship with ours. There was a mutual stimulation, by letters and by discussions during the regular visits which he paid to our country. What follows below is therefore not only a survey of Macdonald’s work but also a piece of history of Dutch mathematics.
Photo credits: First two photos by Peter van Emde Boas, last two photos by Margrit Rösler and Michael Voit.
The roots of Macdonald’s work
For a better understanding of Macdonald’s work one needs to have some notion of a root system. Let $V$ be a Euclidean vector space of dimension $n$ where vectors $\alpha,\beta$ have inner product $(\alpha,\beta)$. For a nonzero vector $\alpha\in V$ we denote by
$s_{\alpha}(\lambda)=\lambda-2\,\frac{(\lambda,\alpha)}{(\alpha,\alpha)}\,\alpha$
the orthogonal reflection of $\lambda$ in $V$ which uses the hyperplane orthogonal to $\alpha$ as a mirror. A root system $R$ in $V$ is a finite set of nonzero vectors in $V$ which spans $V$ and such that
$s_{\alpha}(\beta)\in R\quad\mbox{and}\quad 2\,\frac{(\beta,\alpha)}{(\alpha,\alpha)}\in\mathbb{Z}\quad\mbox{for all \emph{roots} $\alpha,\beta\in R$.}$
Usually it is also required that for any root $\alpha$ the only multiples of $\alpha$ in $R$ are ±$\pm\alpha$. Then $R$ is called a reduced root system. The subgroup $W$ of the orthogonal group $\mathrm{O}(V)$ which is generated by the reflections $s_{\alpha}$ ($\alpha\in R$) is called the Weyl group of $R$.
The theory of root systems and their symmetry groups is a chapter in Euclidean geometry that can be well understood by a second year undergraduate student in mathematics with some knowledge of linear algebra and group theory. The irreducible root systems, i.e., those which cannot be written as an orthogonal union of two smaller ones, are classified by the symbols
$\mathrm{A}_{n},\mathrm{B}_{n},\mathrm{C}_{n},\mathrm{D}_{n},\mathrm{E}_{6},\mathrm{E}_{7},\mathrm{E}_{8},\mathrm{F}_{4},\mathrm{G}_{2}$
and accompanying Coxeter diagrams. By the work of Claude Chevalley in the fifties of the last century we know that they are also the language of the combinatorial infrastructure of the simple algebraic groups $\mathbf{G}(\mathbb{F})$ over a field $\mathbb{F}$. The representation theory of these groups leads to subtle identities in terms of root systems. Conversely such identities make the representation theory computable.
This was first understood around 1925 by Hermann Weyl, who gave an explicit formula, in terms of roots, for the irreducible characters of a compact simple Lie group, considered as a compact real form of $\mathbf{G}(\mathbb{C})$. Weyl proved his formula by transcendental methods. Freudenthal gave in 1954 an algebraic proof of the character formula by means of the quadratic Casimir operator.
Something similar occurred in 1985 in Harish-Chandra’s description [8] of the elementary spherical functions on a Riemannian symmetric space $\mathbf{G}(\mathbb{R})/K$, where $K$ is a maximally compact subgroup of the non-compact simple real Lie group $\mathbf{G}(\mathbb{R})$. These functions are eigenfunctions of a commuting system of linear partial differential operators. Here again the explicit expression, in terms of roots, of the quadratic Casimir operator plays a dominant role.
Important results of Macdonald
In the area just discussed Macdonald has obtained essential results. We present three of them in chronological order.
The first result is the $p$-adic analogue of the elementary spherical functions on $\mathbf{G}(\mathbb{Q}_{p})/\mathbf{G}(\mathbb{Z}_{p})$, announced by Macdonald [21] in 1968. The details appeared in a book [22] which is based on lectures at the Ramanujan Institute in Madras in 1970 and which is written in a very accessible way. One could get a free copy by just sending a request to the author. In a sense, the $p$-adic case is more simple than Harish-Chandra’s real case, because the elementary spherical functions for the $p$-adics are given by an explicit formula (for $\mathbf{G}=\mathbf{GL}_{n}$ in terms of Hall–Littlewood polynomials [26, §V.3]). Macdonald’s explicit formula was one of the ingredients used by Robert Langlands for the formulation of the conjectures named after him [17, p.39].
For a good understanding of the second result we first say somewhat more about Weyl’s character formula. This gives the value of the character on a maximal torus as a quotient of two Fourier polynomials, both skew-symmetric under the Weyl group $W$. The expression in the numerator contains a parameter $\lambda$ (Élie Cartan’s highest weight) which characterizes the irreducible representation. The denominator is the famous Weyl denominator $\Delta$, explicitly given by
$\Delta=\prod_{\alpha>0}(e^{\alpha/2}-e^{-\alpha/2}),$
where the product runs over the set of positive roots $\alpha>0$, i.e., the roots lying on one side of a fixed generic linear subspace of $V$ of codimension 1. When we apply the character formula to the trivial representation, the quotient expresses the trivial character $1$. Hence the numerator specialised at $\lambda=0$ equals the Weyl denominator $\Delta$, which is the celebrated Weyl denominator formula.
In this formula of the form $\prod$=$\sum$ = ∑ the product has terms with opposite signs which cancel each other, by which less terms remain in the sum than first thought. Macdonald observed a similar phenomenon in Jacobi’s triple product identity
$\sum_{m\in\mathbb{Z}}(-1)^{m}q^{m(m-1)/2}z^{m}=\prod_{n=1}^{\infty}(1-q^{n})(1-q^{n}z^{-1})(1-q^{n-1}z).$
This is a classical formula in terms of theta functions in which $q,z$ are complex variables (with $z\neq 0$ and $|q|<1$). In this formula Macdonald recognized the structure of the affine root system of type $\widetilde{\mathrm{A}}_{1}$. Next he found analogous identities for all other affine root systems [23]. By specialization he also obtained simple summation formulas for $\eta^{d}$, where Dedekind’s eta function is given by
$\eta=q^{1/24}\prod_{n=1}^{\infty}(1-q^{n})$
and $d$ is the dimension of a compact simple Lie group. For some values of $d$ this identity was earlier found by the theoretical physicist Freeman Dyson. In an interesting paper [7] he regretted his missed opportunity because the physicist Dyson and the mathematician Dyson had not talked with each other. Alternative proofs of the Macdonald identities were later given by Victor Kac [13] and Eduard Looijenga [20].
Macdonald felt a special sympathy for the general linear group $\mathbf{GL}_{n}$ and the symmetric group $S_{n}$, the “All-embracing Majesties” of the linear algebraic and the finite groups, respectively (as Weyl [40, end of Ch. IV] phrased it for $\mathbf{GL}_{n}$ in ironic adoration). Here the representation theory has a strongly combinatorial flavour, with methods which are, in spite of their elementary nature, not necessarily simple. In 1979 he published the influential book Symmetric Functions and Hall Polynomials on this matter. A much expanded second edition [26] appeared in 1995. For this work he received in 2009 the Leroy P. Steele Prize for Mathematical Exposition.
Macdonald’s contact with the Low Countries
Around 1984 Macdonald visited his good friend Tonny Springer in Utrecht and spoke in the staff colloquium there about his constant term conjectures [24]. The constant term conjecture says that
$\mathrm{CT}(\Delta^{2k})=\prod_{j=1}^{n}\binom{kd_{j}}{k}\quad\mbox{for all $k\in\mathbb{N}$.}$
Here $\Delta$ is the Weyl denominator, $\mathrm{CT}(p)$ stands for the constant term of the Fourier polynomial $p$, and $d_{1},\ldots,d_{n}$ are the degrees of the fundamental invariants of the Weyl group for the root system $R$ of rank $n$ under consideration. For the classical series $\mathrm{A}_{n},\mathrm{B}_{n},\mathrm{C}_{n},\mathrm{D}_{n}$ this formula was known by work of Selberg, Dyson and Good. But for the exceptional root systems $\mathrm{E}_{6},\mathrm{E}_{7},\mathrm{E}_{8},\mathrm{F}_{4},\mathrm{G}_{2}$ it still was an open question.
During the years 1985–87 Gert Heckman and Eric Opdam [9], [10], [31], [32] developed a theory of multivariable hypergeometric functions, depending on a (multi)parameter $k$, which for special values of $k$ coincided with Harish-Chandra’s spherical functions. A crucial step in their work was to show the existence of a system of commuting partial differential operators having as eigenfunctions the functions they searched for. So they were after a natural deformation in $k$ of the commutative algebra of PDOs found by Harish-Chandra in the group case. For symmetric spaces of rank 1, including for instance hyperbolic spaces, the spherical functions are indeed Euler–Gauss hypergeometric functions (in the compact case Jacobi polynomials), as Harish-Chandra (and Élie Cartan in the compact case) already showed. There the extension to general $k$ is immediate. For the root systems $\mathrm{BC}_{2}$ en $\mathrm{A}_{2}$ Tom Koornwinder [14], as part of his thesis in 1975, had already given the commuting differential operators and corresponding polynomial eigenfunctions. Also for general rank these matters had been treated for root systems of type A (Jack polynomials) and BC, including explicit expressions for the commuting PDOs (Sekiguchi, Debiard). But it was the aim of Heckman and Opdam to give a uniform approach for all root systems, not necessarily with explicit expressions.
In January 1988 Opdam defended in Leiden his thesis (essentially the papers [9], [31], [32]), with Macdonald present as a committee member. There, in the final chapter [32], he succeeded to prove by analytic methods the existence of this hypergeometric system for general root systems. This also implied the complete orthogonality of the Jacobi polynomials associated with root systems. Shortly after Opdam showed that, surprisingly, the results of his thesis could be applied to prove the constant term conjecture for general root systems [33].
Almost parallel, in 1987, Macdonald had circulated handwritten manuscripts which described the polynomials now known as Macdonald polynomials. His first manuscript treated a class of symmetric polynomials with two parameters $q$ and $t$. These extrapolated both the Jack polynomials and the Hall–Littlewood polynomials by adding a parameter $q$. He also gave an explicit system of commuting difference operators having his polynomials as eigenfunctions. In the same year Simon Ruijsenaars [37] had independently found this explicit system. Macdonald gives an extensive treatment of this class of symmetric polynomials in [26, Ch. VI]. This chapter also includes the so-called Pieri formula [26, §VI.6] for these polynomials, which Koornwinder [15] proved in 1988 by showing a duality between the variables and the spectral parameters. For this duality Koornwinder was inspired by the two-variable case, where the (homogeneous) Macdonald polynomials are essentially one-variable continuous $q$-ultraspherical polynomials. There he could immediately observe the duality, familiar as he was with the $q$-polynomial stuff by the work of his former teacher Dick Askey [1].
Macdonald’s next two manuscripts in 1987 concerned $q$-analogues of Jacobi polynomials associated with root systems, now known as Macdonald polynomials associated with root systems. They are the direct $q$-analogues of the polynomials introduced by Heckman and Opdam in the same time, which now could be seen as $q\to 1$ limits of Macdonald’s polynomials. For root system of type A Macdonald’s $q$-polynomials are contained in his class of symmetric functions, so it was after the first manuscript a natural step to pass to general root systems. Possibly the work of Heckman en Opdam, which he had closely witnessed by several visits to Leiden in 1987 and 1988, has also been a motivation for him to look at $q$-analogues of their polynomials. The $q$-parameter in Macdonald’s $q$-polynomials gave an additional freedom by which he could prove full orthogonality in an elementary way. And thus, taking the limit for $q\to 1$, the full orthogonality of the Heckman–Opdam polynomials now also had an elementary proof. Macdonald also gave conjectures for the norm and the evaluation at unity of his polynomials. In the limit case $q\to 1$ his conjectures were in accordance with the results of Heckman and Opdam. In 1988 Macdonald put together the two manuscripts into one manuscript [25]. There he also mentioned Opdam’s thesis and proof of the constant term conjecture.
On January 8, 2002, in the aula of the University of Amsterdam, Eric Opdam is reading the laudatio for the honorary doctorate of Macdonald (standing to the right of the pulpit).
Greta and Ian Macdonald in front of attendees of a workshop on the occasion of Tom Koornwinder’s (middle) 60th birthday, all going to embark for a roundtrip through the Amsterdam canals (August 2003).
Further developments in connection with Macdonald polynomials
In the spring of 1989 Heckman [11] gave a toric extension of Charles Dunkl’s work [6] on the differential-reflection operators which are now called Dunkl operators. This enabled him to give an elementary construction of the hypergeometric differential operators which Opdam [32] had constructed by transcendental methods, and it led also to an elementary proof of the orthogonality of their polynomial eigenfunctions, and of the norm and evaluation at unity of the polynomials. A natural question was now how to give a similar construction for Macdonald polynomials. This was quickly answered.
Macdonald during the roundtrip through the canals at the mentioned workshop in Amsterdam, August 2003. Beside him Suzanne Faraut.
In the fall of 1989 the Berlin wall was turned down, and in the summer of the next year many Russians visited the ICM in Kyoto, among whom there was Ivan Cherednik. In a few years all conjectures of Macdonald were made theorems by Cherednik [3]. It was not always easy reading in Ivan’s papers, but expositions such as by Opdam [34] and Macdonald [28] were helpful for understanding. The key to all this were Hecke algebras: first the affine Hecke algebra of Iwahori and Matsumoto coming from the $p$-adic world, next the degenerate Hecke algebra of Drinfeld [5], and for the duality between variable and spectral parameter the double affine Hecke algebra of Cherednik. The commuting generalized Dunkl operators had non-symmetric (non-$W$-invariant) Macdonald polynomials as eigenfunctions, and these were projected onto the symmetric ($W$-invariant) Macdonald polynomials by the symmetrization map.
Another new development were the Koornwinder polynomials [16]. These are a significant 5-parameter extension of the 3-parameter family of Macdonald polynomials associated with the non-reduced root system of type $\mathrm{BC}_{n}$ (here $q$ is not considered as a parameter). This was at the same time an $n$-variable extension of the $4$-parameter family of Askey–Wilson-polynomials [1] in one variable. Also for this case Macdonald formulated conjectures, which again could be proved by Hecke algebra methods [38], [28]. Younger Dutch mathematicians were stimulated by these developments, see van Diejen [4] and Stokman [39] (former students of Ruijsenaars and Koornwinder, respectively). Earlier van der Lek wrote an influential thesis [18]. This was helpful for Ion [12] to give a conceptual proof of the duality theorem for double affine Hecke algebras which was proved in Macdonald’s last book [28, Theorem (3.5.1)] using a case-by-case analysis.
In this period, starting about 1985, the quantum groups also entered the stage. Gradually it became clear that $q=\mbox{quantum}$: Macdonald polynomials and (to a lesser extent) Koornwinder polynomials lived for special parameter values as spherical functions on quantum groups [19][30].
The authors of this paper, each from his own perspective, have experienced the described period as extremely fascinating. Many developments, motivated in quite different ways, interacted and strengthened each other.
Macdonald as a person
We conclude with some words about Ian Macdonald as a person and as a mathematician. He was a very gifted speaker. During his lecture in Séminaire Bourbaki in the middle of the nineties [27] one of us was sitting beside Jim Arthur, an expert in the Langlands program. At the end of the lecture he was silent for a while, and then said: “What an outstanding expositor”. We felt indeed privileged that Macdonald visited the Netherlands so frequently to share his views with us. On these occasions Ian was usually accompanied by his wife Greta, originating from the Flemish part of Belgium. Ian had no driving license, and thus she drove while he read the map for navigation. During a lecture in the early nineties Ian started as follows:
If you have a mathematical problem, there are four things you can do:
1. Ignore it. If you can’t:
2. Prove it. If you can’t:
3. Disprove it. Finally if you can’t do that either:
4. Generalise it.
Everybody was laughing, but of course he was also very serious. If you have already considered options 2 and 3 in vain, then it would be a pity to throw the problem away before having tried option 4. He seemed to apologize that in his lecture he would look again more broadly than asked for by the problem. His conjectured generalizations always turned out to be correct, and in this way he greatly stimulated others.
No doubt Ian was a mathematician with a very classical inclination. During his education in Cambridge the spirit of Hardy and Ramanujan may still have been felt. Ian was a master in dealing with formulas, but in retrospect all these formulas were directed to a better understanding of the representation theory of the simple algebraic groups (over a field or ring). We are truly grateful that the mathematical road travelled by Ian in his later life crossed so frequently with ours.